# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

### v1

@@ Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our strong negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further study the impact of different negative sample construction methods and the number of negative samples on the geometric reasoning performance of LMM, yielding fruitful conclusions. The code and dataset are available at https://github.com/THU-KEG/MMGeoLM.
Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
Kai Sun, Yushi Bai, Zhen Yang, Jiajie Zhang, Ji Qi, Lei Hou, Juanzi Li Tsinghua University
11footnotetext: Equal contribution
1 Introduction
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (LMMs) Anthropic (2024); OpenAI (2023); Bai et al. (2023). It requires fine-grained identification of visual elements Lu et al. (2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them He et al. (2024).
@@ 1 Introduction
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP Patel et al. (2024); Yang et al. (2023); Goel et al. (2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o OpenAI (2024a), Claude-3 Anthropic (2024), and Qwen2.5-VL Bai et al. (2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
To address these shortcomings, recent research has focused strategies such as fine-tuning on specialized mathematical datasets Gao et al. (2023); Zhang et al. (2024c); Peng et al. (2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs Qi et al. (2024); Wang et al. (2024) by aligning images with corresponding captions. However, the construction of image-caption pairs is inherently susceptible to hallucinations, as many captions are generated by existing LMMs. As a result, training with only these positive samples may lead to spurious alignment Zhang et al. (2024b). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Model Training Training Data MM-MATH
@@
AltCLIP MMCLIP ③: Image Negative (4K) 28.2
@@
In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in the vision encoder. For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives. For image-based contrastive learning, we introduce a novel method that leverages a large language model (LLM) to generate a detailed caption and corresponding diagram generation code for a given geometry problem, forming the positive image sample. The LLM then perturbs the code to create visually similar but geometrically incorrect diagrams, which serve as hard negative samples. Additionally, we modify the original CLIP training framework and propose MMCLIP, a method designed to handle an arbitrary number of hard negative samples centered around a single image or caption. Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
We conduct extensive experiments to evaluate the effectiveness of our CLIP training for LMMs. Specifically, we assess our trained MMGeoLM across four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVISTA, and WE-MATH) and open-ended questions (MM-MATH). Our results show that the proposed model outperforms all existing open-source models on GeoQA and MathVISTA, and achieves state-of-the-art performance on MM-MATH, surpassing GPT-4o by 7.5%. We further analyze the impact of different types of hard negative samples. Notably, we find that negative samples constructed from authentic, exam-based images significantly boost model performance—using just 4K image negatives yields better results than over 100K text negatives. Lastly, we observe that increasing the number of hard negatives improves performance up to a certain threshold, beyond which it leads to diminishing returns and eventual degradation.
@@
2 Method
In this section, we first introduce our proposed geometric diagram strong negative construction method in 2.1. Subsequently, we present the corresponding negative captions approach for geometric elements in 2.2. Finally, we describe a training strategy of handling an arbitrary number of negative examples in 2.3, followed by our newly collected multimodal mathematical dataset in 2.4.
2.1 Negative Images Construction
We construct negative geometric examples by generating geometric diagram codes for positive images and perturbing them to create negatives. Existing generation methods Zhang et al. (2024c); Zou et al. (2024); Wei et al. (2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic exam questions, created by educators using professional geometry software McClintock et al. (2002), produce diagrams that are more diverse and accurate. Specifically, in these authentic geometry problems Sun et al. (2024), geometric elements are often described explicitly (e.g., AB is 8 cm) or implicitly through reasoning (e.g., using the Pythagorean theorem). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs OpenAI (2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional informative annotations (see Appendix A for prompt details).
To handle occasional errors in model-generated code, we adopt a lightweight model GLM et al. (2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro DeepMind (2025), avoiding noise from question-based captions. Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, We design prompts (Appendix A) for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
2.2 Negative Captions Construction
We propose two methods for constructing text-based hard negative samples. The first method retrieves captions similar to each positive example from a large image-caption dataset. The second method generates targeted hard negative samples based on reasoning errors observed during LMM inference. Some examples of these hard negative samples are shown in Figure 1.
@@ 2.2 Negative Captions Construction
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance Karpukhin et al. (2020); Xiong et al. (2020); Huang et al. (2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives Xiong et al. (2020); Zhou et al. (2022); Yang et al. (2024). To address this issue, we build upon the recently proposed MAVIS dataset Zhang et al. (2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS Zhou et al. (2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
@@
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships Yuksekgonul et al. (2022); Zhang et al. (2024a), while others leverage in-context learning to generate meaningful modifications Patel et al. (2025). However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level without considering downstream reasoning errors. For example, as shown in Figure 1, LMMs incorrectly interpret the parallel line $AB$ and $CD$ as forming triangle $ABC$, leading to errors in solving the geometric problem. Simply altering positive captions without considering LMM reasoning errors provides limited benefits in improving geometric element recognition.
To address this issue, we analyze the evaluation results of LMMs on MM-MATH Sun et al. (2024) and identify four major types of image element recognition errors in their responses. Based on these error types, we design corresponding rule-based strategies to construct hard negative samples:
@@
    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ are invalid).
@@
We use GLM-4 GLM et al. (2024) to generate 10 rule-based hard negatives for each positive example. An example is shown in Figure 2.
2.3 Hard Negative CLIP Training
Though CLIP-trained vision encoders are widely used in LMMs Radford et al. (2021); Doveh et al. (2023b), they typically adopt in-batch negative sampling during training, which limits their capacity for fine-grained image understanding. While methods like NegCLIP Yuksekgonul et al. (2022) introduce hard negatives, they still rely heavily on in-batch samples. Considering the methods in Section 2.2 have allowed us to construct a large number of hard negatives for each image, we propose MMCLIP training strategy, which focuses on each single image and only uses its corresponding hard negatives for training, entirely abandoning traditional in-batch negatives. In this way, the vision encoder is forced to distinguish subtle differences in all cases, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive example and $T_{j}^{-}$ is the negative samples conducted around the image, using the methods in Section 2.2. Our loss function is formulated as
@@ 2.3 Hard Negative CLIP Training
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings of the input instances. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ are expected to be higher than traditional in-batch negatives. From the above equation, we calculate the gradient for positive and negative samples:
@@
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. We further validate the effectiveness of hard negatives and analyze the impact of their quantity in our experiments in Section 4.2.
2.4 Supervised Fine-tuning Data
@@
3 Multimodal Mathematical Training
3.1 Architecture
We adopt the LLaVA architecture Liu et al. (2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use Mammoth2-7B Yue et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025). The vision encoder is based on AltCLIP Chen et al. (2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
3.2 Training Details
In the vision-text alignment stage, we first pre-train AltCLIP on the 400K MAVIS image-caption alignment dataset Zhang et al. (2024c) to improve the model’s generalization for geometric problems. Next, we fine-tune the model following the MMCLIP training strategy, using 100K hard caption negative samples, each consisting of 10 rule-based and 30 retrieval-based negatives to enhance the model’s understanding of geometric elements. In addition, we fine-tune the model on 4K hard image negative samples, where each example includes 10 negative geometric diagrams. In the second stage, we tune the MLP adapter using 67K G-LLAVA Gao et al. (2023) image-text alignment data. In the third stage, we perform supervised fine-tuning on a combined dataset comprising 300K MAVIS instruction data Zhang et al. (2024c), 117K G-LLaVA instruction data Gao et al. (2023), and 17K open-ended geometric problems. The open-ended dataset includes 12K problems from our collected dataset and 5K randomly sampled MM-MATH Sun et al. (2024) geometric problems, with the rest 700 problems reserved for evaluation. We name the trained model as MMGeoLM.
4 Experiments
To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs. Additionally, we analyze the impact of different vision encoder types and training strategies, as well as the number of hard negatives in MMGeoLM.
4.1 LMM Geometric Reasoning Evaluation
4.1.1 Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA Chen et al. (2021), a geometric question-answering task based on plane geometry; We-Math Qiao et al. (2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista Lu et al. (2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-MATH Sun et al. (2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-MATH, we evaluate models across easy, medium, and hard categories using a test set of 700 problems excluded from training. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
4.1.2 Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus Anthropic (2024), Claude-3.5-Sonnet, GPT-4o OpenAI (2024a), and GPT-4V. Open-source LMMs: MAVIS-7B Zhang et al. (2024c), G-LLaVA-7B Gao et al. (2023), InternVL2-8B Chen et al. (2024), LLaVA-OneVision-7B Li et al. (2024), Qwen2.5-VL-7B-Instruct Bai et al. (2025), Chimera-Reasoner-8B Peng et al. (2024b), InternLM-XComposer2-7B Dong et al. (2024), Phi-3-Vision-128K-Instruct Abdin et al. (2024), and Math-LLaVA-13B Shi et al. (2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks Chen et al. (2021); Lu et al. (2023) since 2023. Additionally, we include human evaluation baselines, using scores from prior studies Zhang et al. (2024c); Sun et al. (2024).
Model GeoQA MathVista We-Math MM-MATH
GPS ALG S2 S3 Easy Med. Hard Avg
@@ 4.1.2 Baselines
Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and MM-MATH. ‘S2/S3’ denote the two-step settings in We-Math. Results marked * are taken from their original papers.
4.1.3 Overall Results

### v2

@@ Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our hard negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations. These analyses yield important insights into optimizing hard negative strategies for geometric reasoning tasks. https://github.com/THU-KEG/MMGeoLM.
Introduction
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic 2024; OpenAI 2023; Bai et al. 2023). It requires fine-grained identification of visual elements (Lu et al. 2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al. 2024).
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al. 2024; Yang et al. 2023; Goel et al. 2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI 2024a), Claude-3 (Anthropic 2024), and Qwen2.5-VL (Bai et al. 2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
@@ Introduction
To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al. 2023; Zhang et al. 2024b; Peng et al. 2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al. 2024; Wang et al. 2024) by aligning images with corresponding captions. However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al. 2024a; Doveh et al. 2023a). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Model Training Training Data $M$-Math
@@
AltCLIP MMCLIP ③: Image-based Negative (4K) 28.2
@@
In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in vision encoder. For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives. For image-based contrastive learning, we introduce a novel method that leverages a large language model (LLM) to generate a detailed caption and corresponding diagram generation code for a given geometry problem, forming the positive image sample. The LLM then perturbs the code to create visually similar but geometrically incorrect diagrams, which serve as hard negative samples. Additionally, we modify the original CLIP training framework and propose MMCLIP, a method designed to handle an arbitrary number of hard negative samples centered around a single image or caption. Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
We evaluate our trained MMGeoLM on four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVista, and We-Math) and open-ended questions (MM-Math). Results show that the proposed model outperforms all existing open-source models on GeoQA and MathVista, and achieves state-of-the-art performance on MM-Math, surpassing GPT-4o by 7.4%. Ablation studies confirm both text-based and image-based hard negatives benefit model performance. Notably, using only 4K image-based negatives yields better results than 100K retrieval-based ones. These findings demonstrate that our hard negative construction and training strategy significantly enhances reasoning accuracy in geometric reasoning.
@@
Method
Negative Images Construction
We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives. Existing generation methods (Zhang et al. 2024b; Zou et al. 2024; Wei et al. 2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic geometry problems (Sun et al. 2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock, Jiang, and July 2002). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al. 2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind 2025). Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
Negative Captions Construction
We propose two methods for constructing text-based hard negative samples. The first method retrieves captions similar to each positive example from a large image-caption dataset. The second method generates targeted hard negative samples based on reasoning errors observed during LMMs’ inference.
@@
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al. 2020; Xiong et al. 2020; Huang et al. 2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al. 2020; Zhou et al. 2022; Yang et al. 2024). To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al. 2024b), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS (Zhou et al. 2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
@@
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al. 2022; Zhang, Awal, and Agrawal 2024), while others leverage in-context learning to generate meaningful modifications (Patel et al. 2025) . However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al. 2024) and identify four major types of image element recognition errors:
@@
    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ is invalid).
@@
To address these error types, we design rule-based strategies that use GLM-4 (GLM et al. 2024) to modify each positive caption and generate 10 hard negative examples.
These text-based negatives are used in model training, and two representative types are shown in Figure 2.
Hard Negative CLIP Training
Figure 3: Overview of the MMGeoLM training pipeline, including main and each ablation experiment configurations, training strategies, and datasets.
Though CLIP-trained vision encoders are widely used in LMMs (Radford et al. 2021; Doveh et al. 2023b), they typically combine random examples for a batch, which limits their capacity for fine-grained image understanding during training. While methods like NegCLIP (Yuksekgonul et al. 2022) introduce hard negatives, they still rely heavily on randomly selected negatives within a batch. Considering the methods in the above section have allowed us to construct a large number of hard negatives, we propose a MMCLIP training strategy, which focuses on each single sample and only uses its corresponding hard negatives for training. In this way, the vision encoder is forced to distinguish subtle differences, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a text-based negative training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive samples and $T_{j}^{-}$ is the negative samples conducted around the image. Our loss function is formulated as follows:
@@
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ is expected to be higher than a randomly chosen one. From the above equation, we calculate the gradient for positive and negative samples:
@@
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. Unlike in-batch training, which relies on batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for rule-based or retrieval-based unimodal negatives. We further validate different training strategies only for hard negatives in the ablation experiments.
Supervised Fine-tuning Data
@@
Multimodal Mathematical Training
Architecture
We adopt the LLaVA architecture (Liu et al. 2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use MAmmoTH2-7B (Yue et al. 2024) and Qwen2.5-7B-Instruct (Qwen et al. 2025). The vision encoder is based on AltCLIP (Chen et al. 2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
Training Details
Model GeoQA MathVista We-Math $M$-Math
GPS ALG S2 S3 Easy Med Hard Avg
@@
Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and $M$-Math. ‘S2/S3’ denotes the two-step settings in We-Math. Results marked * are taken from their original papers.
Our training pipeline, illustrated in Figure 3, consists of three stages covering both main and ablation experiments. In the first stage, we perform vision-text alignment training. For text-based hard negatives, we use the Mavis image-caption alignment dataset (Zhang et al. 2024b). For image-based negatives, we collect 4K geometry questions from real middle-school exams. Additional details, including the types of negatives and the negative-to-positive ratio, are provided in Figure 3. In the second stage, we tune the MLP adapter using 67K image-text pairs from the G-LLaVA dataset (Gao et al. 2023). In the third stage, we conduct supervised fine-tuning. For main experiments, we use a combined dataset comprising 300K Mavis instruction data (Zhang et al. 2024b), 117K G-LLaVA instruction data (Gao et al. 2023), and 17K open-ended geometry questions. For ablation studies, we only use the 17K newly collected geometry problems for efficient validation. We name the final trained model as MMGeoLM.
Experiments
To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs. Additionally, we analyze the impact of different vision encoder types, image-based negatives and training strategies in the subsequent ablation studies.
LMM Geometric Reasoning Evaluation
Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA (Chen et al. 2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al. 2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al. 2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-Math (Sun et al. 2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic 2024), Claude-3.5-Sonnet, GPT-4o (OpenAI 2024a), and GPT-4V. Open-source LMMs: MAVIS-7B (Zhang et al. 2024b), G-LLaVA-7B (Gao et al. 2023), InternVL2-8B (Chen et al. 2024), LLaVA-OneVision-7B (Li et al. 2024), Qwen2.5-VL-7B-Instruct (Bai et al. 2025), Chimera-Reasoner-8B (Peng et al. 2024b), InternLM-XComposer2-7B (Dong et al. 2024), Phi-3-Vision-128K-Instruct (Abdin et al. 2024), and Math-LLaVA-13B (Shi et al. 2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al. 2021; Lu et al. 2023). Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al. 2024b; Sun et al. 2024).
Overall Results